OpenStudy (anonymous):

Find the first partial derivatives of the function: f(x,y,z,t) = xyz^2tan(yt)

OpenStudy (anonymous):

\[f(x,y,z,t) = xyz^2\tan(yt) \]

OpenStudy (unklerhaukus):

\[\frac{\partial}{\partial x} f(x,y,z,t)\\ = \frac{\partial}{\partial x}xyz^2\tan(yt)\\ = \frac{\partial}{\partial x}\big(x\big)\cdot yz^2\tan(yt)+x\cdot \frac{\partial}{\partial x}\big(yz^2\tan(yt)\big)\\=\\=\]

OpenStudy (anonymous):

Partial Derivative in respect to x is: \[yz^2(1)\tan(yt) => yz^2\tan(yt)\]

OpenStudy (unklerhaukus):

yes!

OpenStudy (anonymous):

Cool thanks Now I just have to do the rest of the partial derivatives for y, z and t now. But I'll think I'll take it form here!

OpenStudy (unklerhaukus):

the partial with respect to y is a bit trickier,

OpenStudy (anonymous):

Oh yeah, possibly product rule ~

OpenStudy (unklerhaukus):

yeah, and the chain rule too

OpenStudy (anonymous):

And that too, for that trig function part